Solve the equation: x^2+y^2=34, x+y=2
Answer by ankor@dixie-net.com(22740)

Solve this system of equations:
x^2 + y^2 = 34
x + y = 2
y = (2-x)
:
Replace y in the 1st equation with (2-x)
x^2 + (2-x)^2 = 34
FOIL (2-x)(2-x)
x^2 + 4 - 4x + x^2 = 34
combine like terms
2x^2 - 4x + 4 - 34 = 0
2x^2 - 4x - 30 = 0
Simplify, divide by 2
x^2 - 2x - 15 = 0
Factors to
(x-5)(x+3) = 0
two solutions
x = 5, then y = -3
and
x = -3, the y = +5